A binormal experiment has 6 trials in which p = 0.25. What is the probability of getting at least 3 successes? 

  • 0.8965 

  • 0.0104 

  • 0.1005 

  • 0.1694
Answer #1

p = 0.25

n = 6

P(X > 3) = P(X = 3) + P(X = 4) + P(X = 5) + P(X = 6)

              = 6C3 * 0.253 * 0.753 + 6C4 * 0.254 * 0.752 + 6C5 * 0.255 * 0.751 + 6C6 * 0.256 * 0.750

              = 0.1694

Option-C) 0.1694